You have 15 questions, with 5 answers to each, 1 being the correct choice. What is the probability of getting 2 correct answers?
Answers: .0044
.4288
.4618
.2309

Answer by stanbon(75887)

You have 15 questions, with 5 answers to each, 1 being the correct choice.
What is the probability of getting 2 correct answers?
------
Binomial Problem with n = 15 and p(correct) = 1/5
____
P(x=2) = 15C2*(1/5)^2*(4/5)^13 = binompdf(15,1/5,2) = 0.2309
============================
You either posted the wrong problem or the wrong answers
============
Cheers,
Stan H.
------------

Answers:
.0044
.4288
.4618